solve equation.
x^2-14x=0

Answer by bucky(2189)

Probably the easiest way to solve this equation is by factoring. Since both terms on the left
side contain an x, you can factor out an x to give:
.
x*(x - 14) = 0
.
This equation will be true if either of the factors on the left side are zero because a
multiplication by zero on the left side will make the left side equal the zero on the right side.
.
So we can get the two answers by setting each of the factors equal to zero.
.
For the first factor equal to zero we have:
.
x = 0 and that is one answer.
.
For the second factor equal to zero we have:
.
x - 14 = 0
.
And if we add 14 to both sides we get:
.
x = 14 and that is the second answer.
.
Check ... go to the original equation x^2 - 14x = 0 and substitute 0 for x. The result is:
.
0 - 0 = 0
.
and that is a correct equation, so the answer x = 0 checks.
.
Return to the original equation x^2 - 14x = 0 and substitute 14 for x. When you do you get:
.
14^2 - 14*14 = 0
.
and since 14^2 = 14*14 the equation becomes:
.
14*14 - 14*14 = 0
.
This also is a correct equation so the answer x = 14 checks.
.
Hope this helps you to understand the problem.
.